\begin{thebibliography}{1}
\bibitem{AD} M.B.Abrahamse, R.Douglas, {\em A class of subnormal operators related to multiply connected domains}, Adv. in Math. {\bf 19} (1976), 106–148
\bibitem{Ap} C.Apostol, {\em Functional calculus and invariant subspaces,} J. Operator Theory {\bf 4} (1980), 159-190.
%\bibitem{Clo} R. Clou\^a{}tre,  K R, Davidson, {\em Absolute continuity for row contractions}, J. Functional Analysis {\bf 271}, (2016), 620-641.
%\bibitem{ClT} R. Clou\^a{}tre,E.J.~Timko, {\em Non-commutative measure theory: Henkin and analytic functionals on C*-algebras}, Math. Annalen {\bf 386} (2023), 415-453.
    
% \bibitem{CR} B.~Cole, M.R.~Range, {\em   A-measures on complex manifolds and some applications,} J. Functional Anal. {\bf 11}, (1972), 393-400.

\bibitem{Con} J.~B.~Conway, The Theory of Subnormal Operators, Amer. Math. Soc., Providence Rhode Island 1991.
\bibitem{CrD}
M.J.~Crabb,  A.M.~Davie,   {\em von Neumann’s Inequality for Hilbert Space Operators}, Bull.  London Math. Soc., \textbf{7},  49-50.
\bibitem{E}
J.~Eschmeier, {\em Invariant subspaces for commuting contractions}, J.Operator Theory {\bf 45} (2001), 413-443.

 \bibitem{FS}{C.~Foias, I.~Suciu}, {\em Szeg\"o{} measures and spectral theory in Hilbert spaces}, Rev. Roum. Math, Pures ety Appl. {\bf 11}, (1966) 147-159.

 \bibitem{He} G.M.~Henkin, {\em The Banach spaces of analytic functions in a ball and in a bicylinder are
nonisomorphic,} Funkcional. Anal. i Prilozenia. {\bf 4} (1968), 82–91

 \bibitem{KO}
 M.~Kosiek, A.~Octavio, {\em Representations of $H^\infty(D^N)$, and absolute continuity for N-tuples of contractions},
 Houston J. of Math. {\bf 23}, (1997), 529-537.

\bibitem{MK}
M.~Kosiek, \emph{Absolute continuity of minimal unitary dilations,} Operators and Matrices {\bf 3} (2009), 509-515

\bibitem{WM} W.~Mlak, {\em A note on a Szeg\'o{} type property of semi-spectral measures,} Studia Math. {\bf 31} (1968) 241-251.

\bibitem{mochi}
{N. Mochizuki}, {\em The tensor product of function algebras},{Tohoku Math J.}, {\bf 17} (1965) 139-146.

\bibitem{R}  {K.Rudol},  \emph{The generalised Wold decomposition for subnormal operators,} {
Integral Eqs. Operator Theory} {\bf 11} (1988), 420-436.



 \end{thebibliography}

\documentclass[10pt]{amsart}
\usepackage{graphicx}
\usepackage{amssymb,amsthm,amsmath}
\usepackage{xcolor}
\usepackage{fancyhdr}
\usepackage{hyperref}

\calclayout
\newtheorem{theorem}{Theorem}[section]
\newtheorem{definition}[theorem]{Definition}
\newtheorem{example}[theorem]{Example}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{remark}[theorem]{Remark}

\hypersetup{ colorlinks=true, citecolor=black, linkcolor=black, filecolor=black, urlcolor=black }

\begin{document}
\thispagestyle{empty}
\title[]{Equivalence of Absolute Continuity and Apostol's Condition}
\author[]{Sebastian Foks}
\address{Faculty of Applied Mathematics,  AGH University of Science and Technology,   al. Mickiewicza 30, 30-059,  Krak\'ow, Poland}
\email{sfox@agh.edu.pl }
\keywords{functional calculus, elementary measures, Henkin measure}
\subjclass[2021]{Primary 47A60, 47A13; Secondary 46J15.}

\begin{abstract}
        Absolute continuity of polynomially bounded $n$-tuples of commuting contractions  is studied. A necessary and sufficient condition is found in Constantin Apostol's ``weakened $C_{0,\cdot}$ assumption'', asserting the convergence to 0 of the powers of each operator in a specific topology. Kosiek with Octavio considered  tuples of Hilbert space  contractions satisfying the von-Neumann Inequality. We extend their results to a wider class of tuples, where there may be no unitary dilation and the bounding constant may be greater than 1.  This proof also applies to Banach space contractions and uses decompositions of  A-measures with respect to bands of measures related to Gleason parts of the polydisc algebra.
\end{abstract}

\maketitle

\section{Introduction}
Absolute continuity is one of the important features of operators and it has several meanings depending on the considered type of functional calculus. Perhaps the most frequently used functional calculus is the representation of the algebra $H^\infty(\Omega)$ of bounded analytic functions on a domain $\Omega$. This is related to the invariant subspace problem (still open in the case of Hilbert space operators).  In the case of a single contraction $T$, whose domain is the unit disc, the Nagy-Foias theorem guarantees absolute continuity of completely non-unitary contractions with respect to the Lebesgue measure on the circle. Here, the absolute continuity refers to the spectral measure of the minimal unitary dilation of $T$. This has inspired  the studies of absolute continuity of representations in Hilbert spaces of uniform algebras $A\subset C(X)$ for a compact Hausdorff space $X$.

 For normal operators, one is often interested in properties of their joint spectral measure rather than with these of their dilation (in the contractive case). For a single normal operator $N$, absolute continuity occurs often with respect to the harmonic measure. In particular, if $N$
 is a minimal normal extension of a pure subnormal operator $S$ (i.e. $S$ has no normal part in a reducing, nonzero subspace), it often happens that $\sigma(S)$ is a closure of a domain $\Omega $, while $\sigma(N)$ is  contained in its boundary. Under mild assumptions on  $\Omega$, such absolute continuity allows one to construct a functional model for  $S$ as a bundle shift  (i.e. multiplication by the complex coordinate on a Hardy-type  space of certain analytic sections of a flat unitary bundle over $\Omega$), see \cite{AD},   \cite{R}.

  Absolute continuity of representations of a uniform algebra $A$ with respect to representing measures were first studied with the aid of Szeg\"o{} measures, introduced by Foias with Suciu in their 1966 paper \cite{FS}. These are non-negative Borel measures $\mu$ on $X$ such that for any  subset $E$, with  $\chi_E\cdot L^2(\mu)$ contained in the $L^2(\mu)$-closure $H^2(A,\mu)$ of $A$, one must have $\mu(E)=0$.  A representation $\Phi: A\to \mathcal B(H)$ is called $X$-pure, if the  restrictions  $\Phi(u)|_M$, to a reducing subspace $M\subset H$, cannot form a representation of $A$ in $M$ having an extension to a *-representation of $C(X)$  in $M$, unless $M=\{0\}$. For $X$-pure representations of Dirichlet algebras by \cite[Theorem 7]{FS}, all elementary measures $\mu_f$ for $\Phi$ are Szeg\"o{}.  Mlak has strenghten the absolute continuity results, showing in \cite{WM} that for the  disc algebra, the logarithms of
  Radon-Nikod\'ym derivatives  of Szeg\"o measures with respect to Lebesgue  measure $\lambda $ on $\partial \mathbb D$ are in $L^1(\lambda)$. These results were extended in works of Szyma\'nski, Szafraniec, Kosiek and other members of Mlak's group in Krakow.  

  In the multi-variable case, we consider in the present note   a commuting, polynomially bounded tuple of   contractions. Its absolute continuity   allows one to extend the polynomial functional calculus to a representation of $H^\infty(\mathbb D^n)$ by weak-star continuity.  In this paper, the approach relies on properties of Henkin measures, introduced in 1968 for a different purpose in \cite{He}.

\section{Preliminaries}

Let $\mathcal T :=(T_1, \ldots, T_n)$ be an $n$-tuple of commuting contractions on a Banach space  $H$.   We define the natural representation as $ p(\mathcal T) = p(T_1, \ldots, T_n)$ for
all complex polynomials $p  $ in $n$ commuting variables.
A tuple $\mathcal T$ is 
\emph{polynomially bounded} if, for some constant $C>0$, 
\begin{equation}\label{vNC}
     \| p(T_1, \ldots, T_n) \| \le C \cdot \| p \|_{\infty}, \quad p\in \mathbb C[z_1,\dots z_n],
\end{equation}
where $\| p \|_{\infty}$ is the supremum of $|p|$ over the torus  $\Gamma^n$. 
 The  polynomial functional calculus
  extends then by continuity    to the representation $\Phi^{\mathcal T}$ of the polydisc algebra $A(\mathbb{D}^n)$. This algebra consists of analytic functions on ${\mathbb D^n}$  having continuous extensions to its closure. $A (\mathbb D^n)$ will be considered as a uniform algebra on $\Gamma^n$ (i.e. a closed unital subalgebra of $C(\Gamma^n)$ separating the points of ${\Gamma^n}$).
%
%
  The closed polydisc $\overline{ \mathbb D^n}$ is identified (through evaluation functionals $\delta{}_z$) with the spectrum $\sigma( A(\mathbb{D}^n) )$ of the polydisc algebra.  (Here $\delta{}_z(u):= u(z )$ for $u\in  A(\mathbb{D}^n), \, z\in  \overline{\mathbb{D}^n}$ and $\sigma( A(\mathbb{D}^n) )$ is the set of nonzero  linear and multiplicative functionals $\phi: A(\mathbb{D}^n)\rightarrow \mathbb C$).

   Using the norm in the space $A^*\supset \sigma ( A  )$ of  bounded linear functionals on  a uniform algebra  $A$, we define the equivalence relation $\phi\sim \psi$ on $\sigma (A)$ by
 the inequality $ \| \phi
    - \psi \| < 2 $.

\emph{Gleason parts} of the spectrum of a uniform algebra $A$ are defined as  equivalence classes for the relation $\sim$. For alternative descriptions, see e.g. \cite[Theorem  V.15.8]{Con}.
        By \emph{trivial parts} we mean   parts containing only one point.
For the disc algebra $A(\mathbb D)$, the only parts are $\mathbb D$ and trivial parts $\{\zeta\}$ with $\zeta\in \Gamma$. Explicit description of Gleason parts for the polydisc algebra 
can be deduced  from the fact that $A(\mathbb{D}^n)$ is a tensor product algebra: $A(\mathbb D) \widehat{\otimes}A(\mathbb D^{n-1})$. Its only  Gleason parts are in the form of $n$-fold Cartesian products of Gleason parts for $A(\mathbb D)$. This fact is obtained by applying
 \cite[Theorem 5 and Lemma 3]{mochi} in $n-1$ iterations. It leads to the following observation:
 \begin{proposition}\label{partN}
Any  Gleason part of $\sigma(A(\mathbb D^n))$ that is different from $\mathbb D^n$ is a null-set for $A(\mathbb D^n)$ and is contained in $\{(z_1,\dots, z_n)\in \overline{\mathbb D^n}: z_k=\lambda\}$ for some $k\in\{1,\dots ,n\}$    and $\lambda\in \Gamma$. 
\end{proposition}

\section{Measures on $n$-torus}

  The set $M(X)$ of complex regular Borel measures on a compact Hausdorff space $X$ is a Banach space with total variation norm and $M(X)=C(X)^*$ by Riesz Representation Theorem.

 From now on we consider a polynomially bounded $n$-tuple $\mathcal T$ and the related representation $\Phi^{\mathcal T}: A(\mathbb{D}^n)\rightarrow B(H)$. In the Hilbert space setup, for any pair $x,y\in H$, the inner product $\langle  \Phi^{\mathcal T}(u) x, y \rangle$ defines a bounded linear functional on $A(\mathbb{D}^n)$ with its norm bounded by $ C\|x\|\|y\|$. This functional (non-uniquely)
extends to elements of $C({\Gamma^n})^*$ and there exist  complex Borel measures $\mu_{x,y}\in M({\Gamma^n})$ such that
  \begin{equation}\label{elm}
   \langle \Phi^{\mathcal T}(u) x, y \rangle = \int_{\Gamma^n}  u \ d\mu_{x,y}, \quad  {u \in A(\mathbb{D}^n)}. \end{equation} We call them
  \emph{elementary measures for}  $\Phi^{\mathcal T}$.  In the case of Banach spaces (denoted here also as $H$), one only needs to run $y$ through the dual space $H^*$ and to allow $\langle x, y\rangle$ to denote $y(x)$, the action of a functional  $y$ on a vector $x\in H$, leaving the notation of \eqref{elm} unchanged. Taking $H\times H^*$ as the index set for   families of elementary measures in both cases is hoped to cause no misunderstanding. If $\tilde\mu_{x,y}$ is from another system of elementary measures for the same $\Phi^{\mathcal T}$, then their differences annihilate $A(\mathbb D^n)$. In symbols, $$\mu_{x,y}-\tilde\mu_{x,y} \in A(\mathbb D^n)^\perp:=\{\nu\in M(\Gamma^n): \,\int u\,d\nu=0 \text{ for all } u\in  A(\mathbb D^n)\}.$$

    A \emph{representing measure} at a point $z \in \mathbb{D}^n$ is   a probabilistic Borel measure $\mu$ on $\Gamma^n$ satisfying
$$  \quad u(z) = \int u \ d\mu \quad \text {for all}\quad    {u \in A(\mathbb{D}^n)}$$

If $n=1$, the normalised arc-length is the unique representing measure at 0 for the disc algebra $A(\mathbb D)$ carried on the circle $\Gamma$. However, for $n>1$, there are many such measures on $\Gamma^n$, which leads to the notion of \emph{bands of measures} corresponding to Gleason parts.

In particular, the set $\mathcal M_0 = \mathcal M_0(\Gamma^n )$ of measures $\eta\in M(\Gamma^n)$ absolutely continuous with respect to some representing measure at the point $0$ is \emph{a  band in} $M(\Gamma^n)$, i.e. a closed linear subspace $\mathcal M$ of $M(\Gamma^n)$ that contains all measures  $\eta$ absolutely continuous with respect to some $\mu\in\mathcal M$ (in symbols, $\eta \ll |\mu|$). Measures $\mu_z$ representing points $z\in\mathbb D^n$ belong to  $\mathcal M_0$, but if $z$ belongs to  another Gleason part, then $\mu_z$ is singular to any $\eta\in \mathcal M_0$ \cite{Con}.


\begin{definition}\label{aco}
    A polynomially bounded $n$-tuple  $\mathcal T$ of contractions is \emph{absolutely continuous,} if its natural representation $\Phi^{\mathcal T}$ has a system of elementary measures $\{ \mu_{x,y}
    {(x,y) \in H\times H^*}  \}$  absolutely continuous with respect to certain  measures representing points from $\mathbb{D}^n$  (i.e. if $\mu_{x,y} \in \mathcal M_0$). \end{definition}

 In the case of $n=1$, this amounts to the absolute continuity with respect to the  Lebesgue measure on $\Gamma = \partial \mathbb{D}$. This happens e.g. if $T^n \to 0$ strongly ($T$ is then called a $C_{0,\cdot}$ contraction).

\begin{remark}
If (in the Hilbert space case) $\mathcal T $ has a unitary dilation, then one can ask if the definition above implies absolute continuity of the spectral measure of its minimal unitary dilation. The positive answer is provided in \cite[Theorem 5.1]{MK}.  The existence of such dilation  clearly implies von-Neumann's Inequality, but if $n>2$, either of these properties may fail to hold.
\end{remark}

  Finally, $\mu\in M(\Gamma^n)$ is called an \emph{A-measure} (or a \emph{weak Henkin measure} according to \cite{E}), if  any \emph{Montel sequence} (which means a pointwise convergent to zero on $\mathbb D^n$, bounded sequence of functions $u_k\in A(\mathbb D^n)$) has integrals $\int u_k\,d\mu$ convergent to $0$ as $k\to\infty$. Unlike in the unit ball case, some measures $\nu$ absolutely continuous with respect to a weak Henkin measure  $\mu$ may not share  this property with $\mu$. This leads to the notion of \emph{(strong) Henkin measures}, when any $\nu$ such that $\nu\ll |\mu|$ is also an A-measure. The latter is equivalent to the w-*convergence to 0 in $L^\infty(\mu)$ of any such  $u_k$ (bounded, converging pointwise to 0 on $\mathbb D^n$).

In \cite[Theorem 1.4]{E}, Eschmeier shows that the set of strong Henkin measures  for $  A(\mathbb D^n)$ coincides with the band $\mathcal M_0$. In this proof, the following   result  \cite[Lemma 1.3]{E} is used:
\begin{lemma}\label{EL}
Any weak Henkin measure $\mu\in M(\Gamma^n)$ has a decomposition $ \mu= \mu^a + \mu^s$, where $ \mu^a\in \mathcal M_0, $ and $\mu^s\in M(\Gamma^n)$ is concentrated on a null set of $F_\sigma$-type and $\mu^s$ annihilates $  A(\mathbb D^n)$.
\end{lemma}

\section{Apostol's condition} During his search of  invariant subspaces for a polynomially bounded tuple, Constantin Apostol has formulated in \cite{Ap} the following condition \eqref{Ap3} (weaker than requiring that all $T_j \in C_{0,\cdot}$, as can be seen e.g. when using  absolutely continuous unitary operators):

    \begin{definition}\label{apos} We say that  $\mathcal T= (T_1,\dots, T_n) $ satisfies Apostol's condition  if, for any $(x,y) \in H\times H^*$ and for any $j=1,2,\dots,n$ we have   \begin{equation}\label{Ap3}
                               {\lim_{m\to\infty} \left(\sup_{\| p \|_{\infty} \le 1} | \langle p (T_1, \ldots, T_n)T_j^m x, y \rangle |\right)=0,   ,}\end{equation} where the supremum is taken over all
    polynomials $p \in \mathbb{C}[z_1, \ldots, z_n]$ bounded by 1 on $\mathbb D^n$.
\end{definition}

We shall relate this condition with absolute continuity. In the case when $H$ is a Hilbert space and $\mathcal T$ satisfies von-Neumann's Inequality, which is \eqref{vNC} with $C=1$, the equivalence of these conditions  was obtained by Marek Kosiek with Alfredo Octavio in \cite{KO}. They were  using the so called property (F) which is equivalent to   absolute continuity, but only for $\mathcal T$ satisfying von-Neumann's Inequality.

As mentioned in the remark above, von-Neumann's Inequality is satisfied whenever a unitary dilation exists for $\mathcal T$. This raises the question of whether the result of Kosiek and Octavio can be extended to tuples of Hilbert space operators which lack a unitary dilation and therefore do not always satisfy von-Neumann's Inequality. The following theorem provides a positive answer:

\begin{theorem}\label{MT}
   For  a  polynomially bounded  $n$-tuple of commuting contractions $\mathcal T = (T_1,\dots, T_n)$ in a Banach space $H$, its absolute continuity is  equivalent to
    Apostol's condition \eqref{Ap3}.
\end{theorem}

 \begin{proof}
Assume that \eqref{Ap3} holds. Let us begin  with an arbitrary system of elementary measures $\mu_{x,y}$ for $\Phi^{\mathcal T}$. They are weak Henkin measures. Indeed, let $(u_k)$ be a Montel sequence.  
 From  the proof of \cite[Proposition 1.8]{Ap}, we conclude that $\langle \Phi^\mathcal T(u_k)x,y\rangle =\int u_k d\mu_{x,y} $ converges to zero as $k\to \infty$.  Hence, we may apply Lemma \ref{EL} and decompose each elementary measure as $$\mu_{x,y}= \mu_{x,y}^a + \mu_{x,y}^s, $$ where $\mu_{x,y}^a\in \mathcal M_0$ and $\mu_{x,y}^s \in A(\mathbb{D}^n)^\perp$.  But subtracting an annihilating measure from $\mu_{x,y}$ yields another elementary measure for $\mathcal T$. Thus, we obtain the system $\{\mu_{x,y}^a\}_{(x,y)\in H\times H}$ of elementary measures for the representation $\Phi^\mathcal T$, showing that \eqref{Ap3} implies  absolute continuity.

 To show the opposite implication, assume on the contrary that $\mathcal T$ is absolutely continuous, but \eqref{Ap3} fails. Then for some polynomials $p_j$ bounded by 1 on $\mathbb D^n$  and for some $\epsilon>0$, we would have $|\int p_{j}(z)z_k^j d\mu_{x,y}(z)|>\epsilon$ for infinitely many $j$, despite $p_{j}(z)z_k^j$ forming a bounded sequence in $A(\mathbb D^n)$ and converging point-wise to  zero on $\mathbb D^n$. This contradicts the (strong) Henkin property of $\mu_{x,y}$ valid for any measure from $\mathcal M_0$ by \cite[Theorem 1.4]{E}.
 \end{proof}

 In the following example, we apply the result above for a triple of Hilbert space operators not satisfying von-Neumann's inequality.
 \begin{example}\label{ED} In \cite{CrD}, the following $8\times 8$ (commuting) matrices $T_1,T_2,T_3$ were considered. For  an orthonormal basis $\{e,f_1,f_2, f_3, g_1,g_2,g_3,h\}$  of $\mathbb C^8$, let
 $$T_je=f_j, \,\, T_jf_j=-g_j,\,\,\text{and for } j\neq k \,\, T_jf_k= g_m \text{ where } m\notin\{j,k\},\,\,  T_jg_k=\delta_{jk}h, \,\,T_jh=0. $$ Then it is shown in \cite{CrD} that for $p(z_1,z_2, z_3)=\frac14(z_1z_2z_3 -z_1^3 -z_2^3 -z_3^3)$ one has $\|p\|_\infty <1$ while $\|p(T_1,T_2,T_3)\|\ge 1$ (since $p(T_1,T_2,T_3)e=h$), hence this triple has no unitary dilation.  But on the other hand, each of these operators is nilpotent: an easy calculation shows that $T_j^4=0$. Hence \eqref{Ap3} is satisfied here, as all the powers $T_j^m$ are vanishing for $m > 3$. Also, since  $T_j^2T_k=0$ for any $j\neq k$, the subspace of polynomials annihilating this triple has finite co-dimension in $\mathbb C[z_1,z_2,z_3]$. Indeed, if $n_1+n_2+n_3 >3, n_j\in\mathbb Z_+$, then $T_1^{n_1}T_2^{n_2}T_3^{n_3}=0$.   This easily implies  polynomial boundedness. By Theorem \ref{MT}, this triple is absolutely continuous.
\end{example}


\bibliographystyle{amsplain}

\ifx\undefined\bysame
\newcommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\,}
\fi

\end{document}